OpenStudy (brightemmerald):

What are the vertex and x-intercepts of the graph of the function given below? y = x^2 - 2x - 35 A. Vertex: (0, 0); intercepts: x = -4, -5 B. Vertex: (1, -36); intercepts: x = 7, -5 C. Vertex: (1, -25); intercepts: x = 6, -4 D. Vertex: (7, 5); intercepts: x = 7, 8

OpenStudy (brightemmerald):

@Jaynator495 @zepdrix

OpenStudy (mathmate):

hint: The vertex of a parabola y=ax^2+bx+c is located at x=-b/a therefore the coordinates of the vertex is (-b/a, y(-b/a) )

OpenStudy (brightemmerald):

Do-what?

OpenStudy (brightemmerald):

@ShadowLegendX

OpenStudy (mathmale):

To find the vertex, you need to complete the square of the given function. Have you any experience with completing the square?

OpenStudy (mathmale):

y = x^2 - 2x - 35 can be rewritten as y = x^2 - 2x - 35 Take HALF of the coefficient of x (which is -2); square that and add it to the first two terms, and then subtract the same thing. OK I undrstand you need to go. Come back later.

OpenStudy (mathmale):

Once again: Pls apply all you know about completing the square to finding the vertex here. Do y ou have a printed textbook that might discuss this topic?

OpenStudy (eliesaab):

The x-coordinate of the vertex is \[ -\frac b{2a} \]

OpenStudy (eliesaab):

It is not as mentioned above as -b/a

OpenStudy (eliesaab):

\[ x^2-2x-35=x^2-2x+1-1-35=(x^2-2x +1) -36\] Can you continue from here?

OpenStudy (mathmate):

@eliesaab Thank you for the good catch! Yes, indeed, the coordinates of the vertex should be (-b/(2a), y(-b/(2a) ) However, @mathmale 's suggestion of solving it by completing the square is even better, because there will be nothing to memorize. We see how memory is unreliable! lol